OpenStudy (anonymous):

A sphere has a surface area of 10.1m^2. What is the radius of the sphere to the nearest of a metre? How do i figure this out.

OpenStudy (anonymous):

the equation to find the surface area of a sphere is \[4\pi r^2\]

OpenStudy (anonymous):

so set \[4\pi r^2\]=10.1 and solve for r